OpenStudy (anonymous):

The sum of two numbers is 48. If one third of one number is 5 greater than one sixth of another number, which of the following is the smaller number?

OpenStudy (shamil98):

x + y = 48 1/3 x = 5 + 1/6 y x = 15 + 1/2 y use substitution to solve for y and then for x to get your answer.